In a triangle root(abc(a+b+c))/area is always greater than of equal to ? less than or equal to

In a triangle
root(abc(a+b+c))/area is always greater than of equal to ?
less than or equal to

Grade:9

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ {\sqrt{abc(a+b+c)} \over \Delta}={\sqrt{4R\Delta2s} \over \Delta} \\=>\sqrt{8Rs \over \Delta } \\=>\sqrt{8R \over r} \geq \sqrt{8*2}=4
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty